OpenStudy (anonymous):

what is the equation of the line passig through the points (3,16) , (8,18)

OpenStudy (anonymous):

did you find the slope? you need that first

OpenStudy (dumbcow):

equation of line: \[y - y_1 = m(x - x_1)\] slope \[m = \frac{y_2 - y_1}{x_2 - x_1}\]